### v1

# Synergizing Motion and Appearance: Multi-Scale Compensatory Codebooks for Talking Head Video Generation

Abstract
Talking head video generation aims to generate a realistic talking head video that preserves the person’s identity from a source image and the motion from a driving video. Despite the promising progress made in the field, it remains a challenging and critical problem to generate videos with accurate poses and fine-grained facial details simultaneously. Essentially, facial motion is often highly complex to model precisely, and the one-shot source face image cannot provide sufficient appearance guidance during generation due to dynamic pose changes. To tackle the problem, we propose to jointly learn motion and appearance codebooks and perform multi-scale codebook compensation to effectively refine both the facial motion conditions and appearance features for talking face image decoding. Specifically, the designed multi-scale motion and appearance codebooks are learned simultaneously in a unified framework to store representative global facial motion flow and appearance patterns. Then, we present a novel multi-scale motion and appearance compensation module, which utilizes a transformer-based codebook retrieval strategy to query complementary information from the two codebooks for joint motion and appearance compensation. The entire process produces motion flows of greater flexibility and appearance features with fewer distortions across different scales, resulting in a high-quality talking head video generation framework. Extensive experiments on various benchmarks validate the effectiveness of our approach and demonstrate superior generation results from both qualitative and quantitative perspectives when compared to state-of-the-art competitors.  [^1] https://shaelynz.github.io/synergize-motion-appearance/
1 Introduction
Given a source image and a driving video, talking head video generation (Hong et al., 2022; Tao et al., 2024) aims to animate the person in the source image using the pose and expression from the driving video. Due to its widespread applications, such as video conferencing, the film industry, and virtual reality, it has attracted growing interest in the community.
Significant progress has been made on this task in terms of both quality and robustness in recent years. Existing works primarily focus on learning more accurate motion estimation and appearance representation in 2D or 3D to enhance generation quality. Along the direction, unsupervised methods target predicting local motion flows around unsupervised keypoints without relying on facial priors (Siarohin et al., 2019b; Zhao & Zhang, 2022; Wang et al., 2024), and methods based on predefined models (e.g., 3DMM) (Zakharov et al., 2019; Zhang et al., 2023; Ha et al., 2020) focus on learning robust decoding features to generate high-quality face outputs. Despite the promising achievements, critical challenges persist: 1) Some motion patterns (e.g., local subtle motions) cannot be inferred from a single image pair solely relying on unsupervised keypoints or predefined models for motion estimation, as such models often have limited power of motion representation and may fail to capture certain dynamic aspects of the facial motion from single image pairs (see Fig. 1b). 2) Even with accurate motion estimation, highly dynamic and complex motions in driving videos can create ambiguity during generation, as a still source image lacks sufficient appearance information to handle occluded regions or subtle expression changes (see Fig. 1d). This results in noticeable artifacts and a significant drop in the quality of the generated output. Therefore, generating realistic-looking facial images requires not only inferring accurate motion flow between the two given facial images but also compensating for the intermediate appearance decoding features from the one-shot source image for the final generation of face images.
Figure 1: Effect of motion and appearance compensation with jointly learned compensatory codebooks. Motion flows and warped source appearance features are refined by the complementary information retrieved from the motion and appearance codebooks, which jointly contribute to high-quality generated results.
In this work, we aim to synergize motion and appearance by simultaneously learning accurate motion flows for facial warping and robust facial appearance features for face image decoding, to advance talking head generation. Inspired by the success of codebook learning (Van Den Oord et al., 2017) where compact and useful representations of certain modalities are learned from the whole training dataset, we propose to use such representations as additional knowledge and compensate for motion and appearance with them. Specifically, we propose a unified framework that can achieve joint learning of both motion and appearance codebooks with multi-scale compensation. To refine the motion flow between two facial images (i.e., source and driving), we design a multi-scale motion codebook that captures diverse motion patterns across scales from the entire dataset during training. To enhance intermediate warped facial feature maps for image decoding, we introduce a multi-scale appearance codebook that represents diverse appearance patterns learned from the entire dataset. To use the motion and appearance codebooks, we further introduce a transformer-based compensation structure to iteratively refine motion flows in a coarse to fine manner and refine the warped features across different scales. This approach enables us to improve motion accuracy and capture more facial details by leveraging the diverse motion and appearance information stored in the codebooks. To enhance the learning and compensation of both codebooks, we propose a joint training strategy in which the motion and appearance codebooks are learned simultaneously with the entire framework. This approach allows both codebooks to be optimized together, utilizing gradients from the refined warped features to strengthen their mutual influence and improve overall performance. By learning both multi-scale motion and appearance codebooks, our framework refines the motion flow to accurately warp the source facial features, which are further compensated with additional details from the appearance codebook. This process yields robust intermediate facial decoding features, resulting in improved generation.
We conduct extensive ablation studies to verify the effectiveness of the learned multi-scale motion and appearance codebooks. Experimental results demonstrate that both codebooks effectively enhance the motion flow and intermediate warped features, resulting in more accurate and detailed facial motion flows and feature textures. Furthermore, results on two challenging datasets indicate that our method surpasses state-of-the-art approaches, producing realistic-looking talking head videos. In summary, our contributions are threefold:
  • We propose a novel framework that jointly learns multi-scale motion and appearance codebooks. The motion codebook captures motion patterns at varying levels of granularity, while the appearance codebook stores representative facial structure and texture features. This joint learning enables the model to effectively compensate both motion and appearance for advanced generation.
  • We develop an effective multi-scale compensation mechanism that utilizes the learned motion and appearance codebooks to progressively refine both motion and appearance representations. The mechanism can couple the compensation of both aspects at each level, achieving higher consistency of appearance and motion, thus leading to high visual quality in generated videos.
  • Extensive experiments including on challenging datasets demonstrate that our method not only effectively compensates for facial motions and appearances but also significantly outperforms state-of-the-art approaches, generating more realistic talking head videos.
2 Related Work
Talking Head Video Generation. Existing works on talking head video generation generally separate the motion estimation module and image generation module to disentangle appearance and motion. To transfer the motion, some works require facial priors provided by a pre-trained model during generation. For example, landmark-based approaches (Ha et al., 2020; Zhao et al., 2021; Hong et al., 2024; 2023; Wu et al., 2018; Zakharov et al., 2020) detect pre-defined facial landmarks to transfer the facial pose and expression from a driving frame to the source image. Some other methods (Ren et al., 2021; Zeng et al., 2023; Yao et al., 2020) use parameters from 3D face models (Blanz & Vetter, 2023; Feng et al., 2021; Zhu et al., 2017) as motion descriptors to disentangle identity and pose. However, they normally cannot describe non-facial parts such as hair and neck, and their generation quality is limited by the pre-trained model performance. To address the issue, several methods that do not require any prior knowledge from pre-trained models are proposed. Monkey-Net (Siarohin et al., 2019a) learns sparse motion-related keypoints in an unsupervised manner to describe object movements. FOMM (Siarohin et al., 2019b) extends it with local affine transformation assumption around the keypoints to model complex motion. Subsequent works introduce more flexible mathematical models such as thin-plane spline transformation (Zhao & Zhang, 2022) and continuous piecewise-affine-based transformation (Wang et al., 2024) to increase motion estimation accuracy. Despite the expressiveness of the motion models, they cannot fully describe large head poses and delicate expression changes. MRFA (Tao et al., 2024) tackles the problem by building a correlation volume for each image pair and using it to refine the coarse motion flow iteratively. However, it only uses the warped image feature and a plain image generator for image generation, which may fail when facing extreme pose change as the appearance information from the one-shot source image is not enough. Some works (Yin et al., 2022; Oorloff & Yacoob, 2023; Bounareli et al., 2023) leverage the remarkable generative power of pre-trained StyleGAN (Karras et al., 2019; 2020) for better image generation, but they usually have to balance the editability and fidelity. MCNet (Hong & Xu, 2023) learns a meta-memory bank of spatial facial features to compensate for the warped source features. Different from previous works, we compensate for both motion flows and warped source features with jointly learned multi-scale motion and appearance codebooks to boost the generation quality.
Codebook Learning. Codebook learning aims to learn useful discrete representations with a fixed size. The learned codebook contains rich and compact information, which can facilitate various tasks such as image classification (Cai et al., 2010; Zhang et al., 2009), image synthesis (Esser et al., 2021; Wang et al., 2023b; Chang et al., 2022), blind face restoration (Zhou et al., 2022; Gu et al., 2022) and audio-driven talking head video generation (Wang et al., 2023a). Traditional methods often learn a codebook with unsupervised clustering such as k-means (Csurka et al., 2004). VQ-VAE (Van Den Oord et al., 2017) first incorporates vector quantization in a Variational Autoencoder to learn a codebook containing discrete representative input features. To build a context-rich codebook for images, VQGAN (Esser et al., 2021) further increases the compression rate and adds a discriminator and a perceptual loss on images reconstructed with the codes from the codebook. A transformer is later used to model the composition of the codes for high-resolution image synthesis. CodeFormer (Zhou et al., 2022) uses a learned codebook of compressed high-quality face image features as discrete prior and predicts the code sequence based on the low-quality facial input for blind face restoration. LipFormer (Wang et al., 2023a) learns two codebooks of the upper half face and the bottom half face respectively and predicts the lip codes from the input audio to generate a face video from the audio. We also adopt the idea of codebook learning, but we simultaneously learn multi-scale motion and appearance codebooks that store diverse motion and appearance patterns from the entire dataset during training to facilitate high-quality talking head video generation. The codebooks and the entire framework are trained together so that patterns useful for talking head video generation can be stored in and retrieved from the codebooks.
Figure 2: Overview of the framework. For each scale, multi-scale motion and appearance codebook compensation consists of two sub-modules. (i) Motion Codebook Compensation (MCC) compensates for a motion flow with the motion codebook. (ii) To refine the source facial feature warped by the compensated motion flow, Appearance Codebook Compensation (ACC) produces the compensated appearance feature with the appearance codebook for image decoding. These two sub-modules are employed for all scales. We learn the motion and appearance codebooks jointly with the whole framework.
3 The Proposed Method
In this section, we will present the details of our framework. We jointly learn multi-scale motion and appearance codebooks with compensation for the motion and intermediate appearance features during generation. We adopt the Taylor expansion approximation method to learn the initial motion flow and the warping manner as same as Siarohin et al. (2019a) for video generation.
3.1 Overview
Our framework is illustrated in Fig. 2. First, a keypoint-based motion flow estimator takes both the source image $\mathbf{I}_{s}$ and driving image $\mathbf{I}_{d}$ as input and estimates the initial coarse motion flow $\mathbf{M}^{0}$. An image encoder $\mathcal{E}_{I}$ extracts multi-scale source features $\{\mathbf{F}_{s}^{i}\}_{i=1}^{N}$ from $\mathbf{I}_{s}$. Using $\mathbf{M}^{0}$, $\{\mathbf{F}_{s}^{i}\}_{i=1}^{N}$, and the driving keypoint feature $\mathbf{F}_{kp}^{d}$, the multi-scale motion and appearance codebook compensation module refines the motion flows and warped source features across all scales to obtain the multi-scale compensated appearance features $\{\mathbf{F}_{c}^{i}\}_{i=1}^{N}$ with more accurate motion and less distortion. Finally, the image decoder $\mathcal{D}_{I}$ decodes $\{\mathbf{F}_{c}^{i}\}_{i=1}^{N}$ to generate the final image $\mathbf{I}_{g}$ with the target motion and appearance. Details on the design and learning of multi-scale motion and appearance codebooks and compensation are provided in the following subsections.
3.2 Multi-scale Motion and Appearance Codebook Learning and Compensation
We refine the initial motion flow $\mathbf{M}^{0}$ from coarse to fine and enhance multi-scale source features warped by the refined motion flows with the multi-scale motion and appearance codebook compensation module, producing more accurate motion flows for source feature warping at different scales and restore natural human faces from warping distortions, which together result in higher-quality facial decoding features. Specifically, we jointly learn a multi-scale motion codebook storing local motion flow patterns and a multi-scale appearance codebook containing local facial textures and retrieve relevant information from them. Fig. 3 and Fig. 5 illustrate the motion and appearance codebook learning and compensation process for scale $i$ respectively.
Figure 3: Illustration of motion codebook learning and compensation for scale $i$. We adopt a transformer structure $\mathcal{T}_{M}$ to compensate for the motion flow using the motion codebook. The proposed motion codebook is learned under the supervision of a reconstruction loss and a code-level loss.
3.2.1 Multi-scale Code Allocation
We aim to refine the initial motion flow and the warped source features across all $N$ scales with multi-scale motion and appearance codebooks. As the scale increases, larger appearance features require finer motion flows for accurate warping and more detailed appearance information for appearance compensation. To provide sufficient motion and appearance compensation at larger scales, we introduce a code allocation scheme for multi-scale motion and appearance codebooks, which divides the codebook into multiple groups and allocates more codes for larger scales. We illustrate the scheme with the motion codebook in Fig. 4.
Figure 4: Illustration of the code allocation scheme. We take the multi-scale motion codebook as an example.
The motion codebook $\mathcal{C}_{M}=\{\mathbf{m}_{k}\in\mathbb{R}^{d_{m}}\}_{k=1}^{K}$ contains $K$ codes, and we perform motion compensation at $N$ different scales. The $K$ codes are split into $N$ equal groups, each with $K/N$ codes. At scale $i$, the first $i$ groups, totaling $N_{m}^{i}=i\times K/N$ codes, are allocated for motion compensation. Similarly, for the appearance codebook $\mathcal{C}_{A}=\{\mathbf{a}_{k}\in\mathbb{R}^{d_{a}}\}_{k=1}^{T}$ containing $T$ codes, we also allocate the first $i$ groups, which are the first $N_{a}^{i}=i\times T/N$ codes for appearance compensation. This allows codes with smaller indices to capture general motion patterns or coarse appearance patterns shared across scales, while codes with larger indices to focus on finer motion patterns or facial details needed for larger scales. This maximizes the use of the two codebooks by sharing general information across scales while reserving space for scale-specific details.
At each scale $i$, we form new motion and appearance codebooks $\mathcal{C}_{M}^{i}$ and $\mathcal{C}_{A}^{i}$ from the allocated codes, resulting in $N$ scale-specific motion codebooks $\{\mathcal{C}_{M}^{i}\}_{i=1}^{N}$ and $N$ scale-specific appearance codebooks $\{\mathcal{C}_{A}^{i}\}_{i=1}^{N}$ for multi-scale motion and appearance compensation.
3.2.2 Joint Codebook Learning
To realize effective motion and appearance compensation at different scales, we directly update $\{\mathcal{C}_{M}^{i}\}_{i=1}^{N}$ and $\{\mathcal{C}_{A}^{i}\}_{i=1}^{N}$ with motion flow and appearance feature units at each scale. We jointly optimize the two codebooks with the whole framework, allowing the network to effectively store and retrieve useful patterns at different scales with the codebooks, leading to high-quality talking head video generation.
Motion Codebook Learning. At scale $i$, we use a CNN-based motion flow encoder $\mathcal{E}_{M}$ to map the input motion flow $\mathbf{M}^{i-1}\in\mathbb{R}^{h\times w\times 2}$ into a compact motion flow feature $\mathbf{F}_{m}^{i-1}\in\mathbb{R}^{h_{m}\times w_{m}\times d_{m}}$ where each unit is of length $d_{m}$ and captures a local motion flow pattern on $\mathbf{M}^{i-1}$. We quantize each of its spatial element $\mathbf{F}_{m}^{i-1}(x,y)$ with its nearest code from $\mathcal{C}_{M}^{i}$ to obtain a quantized motion flow feature $\hat{\mathbf{F}}_{m}^{i-1}$:
$\hat{\mathbf{F}}_{m}^{i-1}=\mathit{Q}(\mathbf{F}_{m}^{i-1}):=\left(\mathop{\arg\min}_{\mathbf{m}_{k}\in\mathcal{C}_{M}^{i}}||\mathbf{F}_{m}^{i-1}(x,y)-\mathbf{m}_{k}||_{2}^{2}\right).$ (1)
A motion flow decoder $\mathcal{D}_{M}$ reconstructs $\mathbf{M}^{i-1}$ with the quantized motion flow feature $\hat{\mathbf{F}}_{m}^{i-1}$:
$\hat{\mathbf{M}}^{i-1}=\mathcal{D}_{M}(\hat{\mathbf{F}}_{m}^{i-1})=\mathcal{D}_{M}(\mathit{Q}(\mathcal{E}_{M}(\mathbf{M}^{i-1}))).$ (2)
To update the scale-specific motion codebook $\mathcal{C}_{M}^{i}$ with local motion flow patterns from $\mathbf{F}_{m}^{i-1}$, we use the following loss function:
$\begin{split}\mathcal{L}_{vq,m}^{i}=\lambda_{recon,m}||\hat{\mathbf{M}}^{i-1}-sg[\mathbf{M}^{i-1}]||_{1}&+||sg[E_{M}({\mathbf{M}}^{i-1})]-\hat{\mathbf{F}}_{m}^{i-1}||_{2}^{2}\\ &+\beta||sg[\hat{\mathbf{F}}_{m}^{i-1}]-E_{M}(sg[\mathbf{M}^{i-1}])||_{2}^{2},\end{split}$ (3)
where $sg[\cdot]$ denotes the stop gradient operator, and $\lambda_{recon,m}$ and $\beta$ are the loss term weights. The first term represents the motion flow reconstruction loss, while the last two terms form a code-level loss (Van Den Oord et al., 2017) that minimizes the distance between the latent motion flow units and the motion codes. We stop the gradient of $\mathbf{M}^{i-1}$ to ensure that motion codebook learning at scale $i$ does not interfere with the training of the motion flow estimator or the compensated motion flows from prior scales. The overall loss function for motion codebook learning across all $N$ scales is $\mathcal{L}_{vq,m}=\sum_{i=1}^{N}\mathcal{L}_{vq,m}^{i}$.
Appearance Codebook Learning. We use the image encoder $\mathcal{E}_{I}$ to extract multi-scale driving features $\{\mathbf{F}_{d}^{i}\}_{i=1}^{N}$, which serve as targets for our appearance codebook. Since image features of different scales have varying resolutions, directly flattening high-resolution features for compensation can be computationally expensive. To address this, we apply window partitioning. For a feature map of shape $(h_{a}^{i},w_{a}^{i},c_{a}^{i})$ at scale $i$, we divide it into patches of shape $(h_{a}^{i}/h_{a},w_{a}^{i}/w_{a})$, reshaping the feature into $(h_{a},w_{a},c_{a}^{i}\times h_{a}^{i}\times w_{a}^{i}/h_{a}/w_{a})$. We then linearly project the features into $d_{a}$ dimensions to align them with the appearance codes. This results in a more compact driving feature $\overline{\mathbf{F}}_{d}^{i}\in\mathbb{R}^{h_{a}\times w_{a}\times d_{a}}$, where each unit represents an appearance pattern at scale $i$. Finally, element-wise quantization with the nearest code from $\mathcal{C}_{A}^{i}$ produces the quantized appearance feature $\overline{\mathbf{F}}_{d}^{i\prime}$:
$\overline{\mathbf{F}}_{d}^{i\prime}=\mathit{Q}(\overline{\mathbf{F}}_{d}^{i}):=\left(\mathop{\arg\min}_{\mathbf{a}_{k}\in\mathcal{C}_{A}^{i}}||\overline{\mathbf{F}}_{d}^{i}(x,y)-\mathbf{a}_{k}||_{2}^{2}\right).$ (4)
To update the scale-specific appearance codebook $\mathcal{C}_{A}^{i}$ with local appearance units from $\overline{\mathbf{F}}_{d}^{i}$, we use the following loss function:
$\mathcal{L}_{vq,a}^{i}=||sg[\overline{\mathbf{F}}_{d}^{i}]-\overline{\mathbf{F}}_{d}^{i\prime}||_{2}^{2}+\beta||sg[\overline{\mathbf{F}}_{d}^{i\prime}]-\overline{\mathbf{F}}_{d}^{i}||_{2}^{2},$ (5)
where $sg[\cdot]$ denotes the stop gradient operator, and $\beta$ is the loss term weight. We only use the code-level loss to reduce the distance between the appearance feature units and the codes from $\mathcal{C}_{A}^{i}$. The overall loss function for appearance codebook learning is $\mathcal{L}_{vq,a}=\sum_{i=1}^{N}\mathcal{L}_{vq,a}^{i}$. We do not use the image decoder $\mathcal{D}_{I}$ to reconstruct the original driving image $\mathbf{I}_{d}$ from $\{\overline{\mathbf{F}}_{d}^{i\prime}\}_{i=1}^{N}$, allowing $\mathcal{D}_{I}$ to focus on decoding the compensated appearance features and improving the talking head video generation. Training a separate image decoder for reconstruction with quantized appearance codes would be computationally expensive. Experimental results in Sec. 4.3 show that using only the code-level loss can already learn the appearance codebook effectively.
Figure 5: Illustration of appearance codebook learning and compensation for scale $i$. We utilize a transformer structure $\mathcal{T}_{A}$ to compensate for the warped features with the appearance codebook, adding more facial details to the feature map. Similar to the motion codebook, the designed appearance codebook is learned under the supervision of a code-level loss.
3.2.3 Multi-scale Joint Codebook Compensation
Motion and Appearance Codebook Compensation Coupling. With the multi-scale motion and appearance codebooks, we couple motion and appearance codebook compensation at each scale to produce better image decoding features, as shown in Fig. 2. Specifically, at each scale, motion codebook compensation (MCC) refines the input motion flow $\mathbf{M}^{i-1}$ to obtain the compensated motion flow $\mathbf{M}^{i}$, which is used to warp the source feature $\mathbf{F}_{s}^{i}$. Appearance codebook compensation (ACC) then refines the warped source feature to produce the compensated appearance feature $\mathbf{F}_{c}^{i}$. If $i<N$, $\mathbf{M}^{i}$ is used as the input motion flow for the next scale. In this way, motion and appearance codebook compensation are coupled across scales, enabling more consistent motion and appearance for high-quality video generation.
Motion Codebook Compensation. An intuitive approach for multi-scale motion codebook compensation is to retrieve motion codes from the scale-specific codebook and decode them into finer motion flows using $\mathcal{D}_{M}$ at each scale. However, to reduce computational complexity, we limit the number of motion codes, which decreases expressiveness. This makes it difficult to fully reconstruct motion flows, leading to degraded image quality. Instead, we predict motion flow residual for the input flow at each scale, allowing the motion codebook to enhance the motion flow without requiring precise reconstruction.
As shown in Fig. 3, to retrieve motion residuals from $\mathcal{C}_{M}^{i}$ at scale $i$, we use a motion code retrieval transformer $\mathcal{T}_{M}$ shared for all scales. It queries the scale-specific codebook $\mathcal{C}_{M}^{i}$ using the encoded motion feature $\mathbf{F}_{m}^{i-1}$, the warped source feature $\mathbf{F}_{cw}^{i-1}$, and the driving keypoint feature $\mathbf{F}_{kp}^{d}$. The first two represent the current motion flow, and the last indicates the target pose. These features are processed through a convolutional encoding block and concatenated into a compact motion query feature, then flattened and enhanced with a learnable position embedding before being passed to $\mathcal{T}_{M}$. $\mathcal{T}_{M}$ consists of $L_{M}$ transformer layers, each with multi-head self-attention, cross-attention, and convolution layers (instead of linear layers) to preserve spatial structure. The self-attention models global correlations, and the cross-attention uses the output of self-attention as the query and the codes from $\mathcal{C}_{M}^{i}$ as key-value pairs to retrieve local motion flow patterns. The transformer outputs the motion flow residual feature $\mathbf{F}_{mr}^{i-1}$, which is decoded by a motion flow residual decoder $\mathcal{D}_{MR}$ to obtain the motion flow residual $\mathbf{M}_{r}^{i}\in\mathbb{R}^{h\times w\times 2}$. Finally, we add $\mathbf{M}_{r}^{i}$ to the input motion flow $\mathbf{M}^{i-1}$ to produce the compensated motion flow $\mathbf{M}^{i}$, which is used for source feature warping at scale $i$.
The compensated motion flow $\mathbf{M}^{i}$ is sufficient for warping the source feature $\mathbf{F}_{s}^{i}$, but may lack fine details for higher-resolution features like $\mathbf{F}_{s}^{i+1}$. Therefore, we use $\mathbf{M}^{i}$ as input for motion codebook compensation at scale $i+1$ and refine it with more motion codes. This iterative process continues across scales, producing multi-scale compensated motion flows $\{\mathbf{M}^{i}\}_{i=1}^{N}$.
Appearance Codebook Compensation. At scale $i$, we first warp the source feature $\mathbf{F}_{s}^{i}$ with the compensated motion flow $\mathbf{M}^{i}$ to obtain the warped feature $\mathbf{F}_{w}^{i}$. We then apply window partitioning to map $\mathbf{F}_{w}^{i}$ into a compact feature $\overline{\mathbf{F}}_{w}^{i}$. To correct the corrupted appearance in $\overline{\mathbf{F}}_{w}^{i}$, we retrieve appearance codes with $\overline{\mathbf{F}}_{w}^{i}$ using a transformer $\mathcal{T}_{A}$. Similar to motion codebook compensation, $\overline{\mathbf{F}}_{w}^{i}$, reshaped and augmented with a learnable position embedding, passes through $L_{A}$ transformer layers where appearance codes from $\mathcal{C}_{A}^{i}$ are retrieved with cross-attention. The transformer output is the compensated appearance feature $\overline{\mathbf{F}}_{c}^{i}$, which retains the pose but reduces distortion of $\overline{\mathbf{F}}_{w}^{i}$. Finally, we apply window reverse on $\overline{\mathbf{F}}_{c}^{i}$ to restore it to the original image shape $(h_{a}^{i},w_{a}^{i},c_{a}^{i})$ using linear projection and reshaping. This compensation is performed across all scales, resulting in multi-scale compensated appearance features $\{\mathbf{F}_{c}^{i}\}_{i=1}^{N}$.
3.3 Joint Optimization Objective of the Framework
We use the multi-scale compensated appearance features $\{\mathbf{F}_{c}^{i}\}_{i=1}^{N}$ and the image decoder $\mathcal{D}_{I}$ for image decoding. The low-resolution feature $\mathbf{F}_{c}^{1}$ is fed as the initial input to $\mathcal{D}_{I}$, which gradually upsamples it through a series of upsampling layers and ResNet blocks (He et al., 2016) until it reaches the output resolution. The intermediate features are refined with higher-resolution features $\{\mathbf{F}_{c}^{i}\}_{i=2}^{N}$ by SFT (Wang et al., 2018) and addition when they match the resolution of $\mathbf{F}_{c}^{i}$. After fusing all the scales, $\mathcal{D}_{I}$ generates the final output $\mathbf{I}_{g}$.
The training objective combines the codebook losses from Sec. 3.2 with common losses for talking head video generation (Siarohin et al., 2019b; Hong et al., 2022). Specifically, we use the equivariance loss $\mathcal{L}_{eq}$ and keypoint distance loss $\mathcal{L}_{kpd}$ to guide keypoint prediction in the motion flow estimator, and an image reconstruction loss $\mathcal{L}_{recon}$ and an adversarial loss $\mathcal{L}_{adv}$ on $\mathbf{I}_{g}$. To avoid the image decoder relying too much on high-resolution appearance features, we also generate an image $\mathbf{I}_{g}^{1}$ using only $\mathbf{F}_{c}^{1}$ and apply $\mathcal{L}_{recon}$ to minimize the difference between $\mathbf{I}_{g}^{1}$ and $\mathbf{I}_{d}$. The overall training objective is:
$\mathcal{L}=\mathcal{L}_{recon}(\mathbf{I}_{d},\mathbf{I}_{g})+\lambda_{adv}\mathcal{L}_{adv}(\mathbf{I}_{d},\mathbf{I}_{g})+\mathcal{L}_{eq}+\mathcal{L}_{kpd}+\mathcal{L}_{vq,m}+\mathcal{L}_{vq,a}+\lambda^{1}\mathcal{L}_{recon}(\mathbf{I}_{d},\mathbf{I}_{g}^{1}),$ (6)
where $\lambda_{adv}$ and $\lambda^{1}$ are the corresponding loss term weights.
4 Experiments
4.1 Implementation Details
Datasets. We conduct experiments on VoxCeleb1 (Nagrani et al., 2017) and CelebV-HQ (Zhu et al., 2022) datasets. We train our model on VoxCeleb1 training set. For evaluation, we build the test set on VoxCeleb1 by randomly sample 50 videos from its test split. To evaluate the model’s generalization ability, we randomly select 50 videos from CelebV-HQ for testing.
Metrics. For same-identity reconstruction, we adopt PSNR, $\mathcal{L}_{1}$ and LPIPS following (Tao et al., 2024) to evaluate the reconstruction quality. We also use FID (Heusel et al., 2017) to measure the realism of the generated video frames. Following (Siarohin et al., 2019a), we employ Average Keypoint Distance (AKD) for motion transfer quality evaluation and Average Euclidean Distance (AED) for identity preservation quality evaluation.
4.2 Comparison with State-of-the-art Methods
Figure 6: Qualitative comparison with state-of-the-art methods for (a) same-identity reconstruction and (b) cross-identity reenactment on the VoxCeleb1 and CelebV-HQ datasets. Our method can generate high-quality facial images under even large pose variations.
Table 1: Quantitative comparison with state-of-the-art methods for same-identity reconstruction on VoxCeleb1 and CelebV-HQ dataset. Our results are the best on the Voxceleb1 dataset and competitive on the CelebV-HQ dataset.
Method VoxCeleb1 CelebV-HQ
FID $\downarrow$ PSNR $\uparrow$ $\mathcal{L}_{1}$downarrow LPIPS $\downarrow$ AKD $\downarrow$ AED $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ $\mathcal{L}_{1}$downarrow LPIPS $\downarrow$ AKD $\downarrow$ AED $\downarrow$
FOMM (Siarohin et al., 2019b) 53.97 22.96 0.0474 0.2200 1.4037 0.1509 78.15 20.92 0.0685 0.2925 3.6098 0.2955
DaGAN (Hong et al., 2022) 51.55 22.92 0.0492 0.2251 1.5740 0.1652 99.84 20.49 0.0781 0.3209 7.6075 0.3301
MCNet (Hong & Xu, 2023) 51.45 24.59 0.0402 0.1996 1.2363 0.1254 78.33 22.20 0.0640 0.2732 4.1386 0.2903
MRFA (Tao et al., 2024) 48.49 25.26 0.0370 0.1872 1.1823 0.1188 75.73 22.41 0.0625 0.2670 3.7166 0.2527
AniPortrait (Wei et al., 2024) 52.65 20.15 0.0637 0.2767 2.6543 0.2623 61.56 19.71 0.0748 0.2878 2.2397 0.2739
FYE (Ma et al., 2024) 43.25 19.54 0.0714 0.2954 2.7071 0.2652 62.55 19.58 0.0802 0.3006 4.8637 0.3029
Ours 43.15 25.30 0.0355 0.1846 1.2039 0.1071 71.78 22.40 0.0610 0.2608 3.2562 0.2825
We compare our method with a series of open-source state-of-the-art methods including non-diffusion based FOMM (Siarohin et al., 2019b), DaGAN (Hong et al., 2022), MCNet (Hong & Xu, 2023) and MRFA (Tao et al., 2024), and diffusion-based AniPortrait (Wei et al., 2024) and Follow-Your-Emoji (FYE) (Ma et al., 2024).
Same-identity Reconstruction. To evaluate the performance on same-identity reconstruction, we use the first frame of each video as the source image and reconstruct the whole video. Tab. 1 presents the quantitative results for same-identity reconstruction. Our method outperforms the other methods almost on all metrics on VoxCeleb1 dataset and remains competitive when generalized to the more challenging CelebV-HQ. Compared with those unsupervised methods (Siarohin et al., 2019b; Hong et al., 2022; Hong & Xu, 2023; Tao et al., 2024), our method achieves better motion estimation, e.g., our method get the best AKD scores on CelebV-HQ dataset. This result verifies the effectiveness of our designed multi-scale motion codebook and its generalizability. For the results of image quality (i.e., FID, PSNR, $\mathcal{L}_{1}$, LPIPS), our method outperforms other methods on VoxCeleb1 dataset, even those diffusion methods (Wei et al., 2024; Ma et al., 2024) trained with larger-scale datasets. It indicates that our designed multi-scale appearance codebook is capable to compensate for the intermediate warped feature for better talking head video generation.We also show some qualitative results in Fig. 6a. Our method can generate plausible unseen facial regions (row 2) and handle large motion (row 4) with accuracy.
Cross-identity Reenactment. We conduct cross-identity reenactment experiments to validate our method. The qualitative results shown in Fig. 6b indicate the superiority of our method. Compared to non-diffusion-based MCNet (Hong & Xu, 2023) and MRFA (Tao et al., 2024), our method better preserves facial details (row 1 and 3), even under large motions (row 2 and 4). Diffusion-based FYE (Ma et al., 2024) often produced exaggerated expressions and struggled to imitate the driving expressions accurately, likely due to its reliance on landmark-based embeddings without explicitly modeling motion. These findings confirm the effectiveness of our framework. More experimental results are shown in Sec. A.2.2.
4.3 Ablation Study
We perform ablation studies to assess the effectiveness of the learned multi-scale motion and appearance compensatory codebooks. The model variants in Tab. 2 are as follows: (i) “Baseline” is the model without any compensatory codebook. (ii) “Baseline+SMC” includes only a single-scale motion codebook, compensating the initial motion flow only at scale $1$, and the compensated result is used to warp multi-scale features. (iii) “Baseline+MMC” includes a multi-scale motion codebook to compensate for the motion flows across scales. (iv) “Baseline+MMC+SAC” adds a single-scale appearance codebook to (iii), compensating only the warped feature at scale $1$. (v) “Baseline+MMC+MAC” is our full model with both multi-scale motion and appearance codebooks. We present the quantitative results in Tab. 2 and qualitative comparisons of (i), (iii), and (v) in Fig. 9
Effect of Joint Codebook Learning. To evaluate the effectiveness of our jointly learned multi-scale motion and appearance codebooks, we visualize the reconstructed multi-scale motion flows with the motion codebook in Fig. 7 and appearance features with the appearance codebook in Fig. 8. For the motion flows, results in Fig. 7 are output of the motion flow decoder $\mathcal{D}_{M}$ given the quantized motion flow features. Despite the limited number of codes, our multi-scale motion codebook can reconstruct high-quality motion flows, confirming its ability to capture typical local motion patterns. For appearance features, we visualize the quantized features directly in Fig. 8. Despite some quantization loss, the multi-scale appearance codebook reconstructs the driving features well with limited codes, demonstrating the its ability to store informative local appearance details. These results demonstrate the effectiveness of joint codebook learning, allowing the codebooks to store expressive local motion and appearance patterns.
Figure 7: Visulization of the original and reconstructed motion flows for different scales. Our multi-scale motion codebook can reconstruct multi-scale motion flows with high quality.
Figure 8: Visulization of the original and reconstructed driving features. Our multi-scale appearance codebook can reconstruct multi-scale appearance features with acceptable quantization loss.
Table 2: Ablation study on the multi-scale motion and appearance codebook compensation. We present the results for same-identity reconstruction on VoxCeleb1 dataset.
FID $\downarrow$ PSNR $\uparrow$ $\mathcal{L}_{1}$downarrow LPIPS $\downarrow$ AKD $\downarrow$ AED $\downarrow$
Baseline 47.83 24.93 0.0375 0.1954 1.2384 0.1106
+ SMC 49.00 24.97 0.0371 0.1917 1.2183 0.1167
+ MMC 44.86 25.27 0.0360 0.1875 1.2171 0.1076
+ MMC + SAC 44.32 25.16 0.0362 0.1856 1.2106 0.1091
+ MMC + MAC (Ours) 43.15 25.30 0.0355 0.1846 1.2039 0.1071
Figure 9: Qualitative ablation study on multi-scale motion and appearance codebook compensation. Both motion and appearance codebook compensation contribute to better generation quality.
Effect of Multi-scale Motion Codebook Compensation. The second row of Tab. 2 shows that using single-scale motion codebook compensation already improves motion transfer and image quality, reflected by better PSNR, $\mathcal{L}_{1}$, LPIPS, and AKD compared to the baseline. This highlights the effectiveness of motion codebook compensation for talking head video generation. Multi-scale motion codebook compensation further enhances all metrics, underscoring the importance of handling motion flows at different scales for improved feature warping. Fig. 9 shows that multi-scale motion compensation achieves better motion transfer (e.g., head pose in row 1), reduces artifacts (e.g., hair in row 2), and preserves source identity (row 3). Fig. 10 visualizes the motion flow compensation process, showing that the initial motion flow $\mathbf{M}^{0}$ is rough, but multi-scale compensation refines it iteratively with residuals $\{\mathbf{M}_{r}^{i}\}_{i=1}^{N}$, adding finer details as the scale increases for smoother, more face-adapted motion flows.
Figure 10: Visualization of the motion flow compensation process. We present the initial motion flow $\mathbf{M}^{0}$, the motion flow residual $\{\mathbf{M}_{r}^{i}\}_{i=1}^{N}$ and the compensated motion flows $\{\mathbf{M}^{i}\}_{i=1}^{N}$.
Figure 11: Visualization of the appearance compensation results. We present the warped source features $\{\mathbf{F}_{w}^{i}\}_{i=1}^{N}$ and the compensated appearance features $\{\mathbf{F}_{c}^{i}\}_{i=1}^{N}$.
Effect of Multi-scale Appearance Codebook Compensation. The fourth row in Tab. 2 shows that adding single-scale appearance codebook compensation on top of multi-scale motion codebook compensation improves FID, LPIPS, and AKD, indicating that appearance codebook compensation refines warped source features for more realistic image generation. However, there is a slight drop in PSNR, $\mathcal{L}_{1}$, and AED, likely due to a conflict between scale $1$ compensated appearance features and other warped features with warping artifacts. The fifth row shows consistent improvement, suggesting multi-scale appearance codebook compensation resolves this conflict and further refines warped features across scales, boosting overall performance. The last two columns in Fig. 9 also verify that multi-scale appearance codebook compensation leads to more accurate motion (e.g., the mouth in row 1, eyes in row 2, and shoulders in row 3) with realistic facial details (e.g., the hair in row 2). Additionally, Fig. 11 visualizes the compensated feature maps at different scales, showing more complete facial shapes and details compared to the warped feature $\mathbf{F}_{w}^{i}$. These results validate the effectiveness of multi-scale appearance codebook compensation.
5 Conclusion
In this paper, we have presented a novel framework that jointly learns multi-scale motion and appearance compensatory codebooks to enhance the motion flows and appearance features for talking head video generation. The motion and appearance codebooks store local motion and appearance patterns learned from the entire dataset at different scales, and our multi-scale motion and appearance codebook compensation module retrieves useful codes from the codebooks with a transformer-based strategy at different scales to refine motion flows and appearance features for image generation. Extensive results verify the effectiveness of our joint motion and appearance codebook learning and compensation, synergizing both for high-quality talking video generation.
Appendix A Appendix
A.1 Additional Implementation Details
We perform multi-scale compensation across $N=4$ scales. We employ the keypoint-based motion flow estimator from FOMM (Siarohin et al., 2019b). The multi-scale motion flows are estimated at a size of $64\times 64$. We use convolution layers to encode the motion flows into a latent motion flow space of size $32\times 32\times 32$ and set the multi-scale motion codebook size to $K=1024$ and $d_{m}=32$. We also use convolution layers to decode the quantized motion flow features while adopting the motion flow updater in MRFA (Tao et al., 2024) as our motion flow residual decoder. We employ the image encoder and decoder architecture from VQGAN (Esser et al., 2021) and further encode the multi-scale appearance features into a size of $32\times 32\times 256$. The multi-scale appearance codebook size is set to $T=1024$ and $d_{a}=256$.
We follow the unsupervised training pipeline from FOMM (Siarohin et al., 2019b), where the source and driving frames are extracted from the same video, and our framework learns to reconstruct the driving frame. For the training objective, we use the perceptual loss from FOMM (Siarohin et al., 2019b) along with the L1 loss as the image reconstruction loss, and we set the loss weights as $\lambda_{adv}=0.8$, $\lambda^{1}=0.5$, $\lambda_{recon,m}=32$ and $\beta=0.25$. We train the entire framework end-to-end using the Adam optimizer with a learning rate of $8\times 10^{-5}$ and a batch size of 16 for 250K iterations on four NVIDIA RTX 3090 GPUs.
A.2 More Details on Experiments
A.2.1 Additional Details on the Compared Methods
We evaluate the performance of the compared methods using their released pre-trained models, and we present the training datasets used for each method in Tab. 3. All the GAN-based methods (Siarohin et al., 2019b; Hong et al., 2022; Hong & Xu, 2023; Tao et al., 2024) and our method are trained on the VoxCeleb1 (Nagrani et al., 2017) training set, while the diffusion-based methods AniPortrait (Wei et al., 2024) and Follow-Your-Emoji (FYE) (Ma et al., 2024) are trained on larger-scale datasets, including VFHQ (Xie et al., 2022), CelebV-HQ (Zhu et al., 2022), HDTF (Zhang et al., 2021), and their own collected dataset (Ma et al., 2024).
A.2.2 More Experimental Results
Video Demo. We present a video demo along with the supplementary material, which includes video results from the ablation study and comparisons with state-of-the-art methods to demonstrate the performance of our video generation method.
Figure 12: User study results ranking the quality of videos generated by different methods.
Cross-identity Reenactment. Since there is no ground truth for cross-identity reenactment, we conduct a user study comparing our approach to recent state-of-the-art methods, including a GAN-based model (MRFA (Tao et al., 2024)) and a diffusion-based model (Follow-You-Emoji (FYE) (Ma et al., 2024)). We randomly selected 10 source-driving pairs and asked 30 participants to evaluate the generated videos based on appearance realism, motion naturalness, and overall quality. The results shown in Fig. 12 indicate that users prefer our method, confirming its superiority.
Table 3: Quantitative comparison with state-of-the-art methods for cross-identity reenactment on VoxCeleb1 dataset. AniPortrait (Wei et al., 2024) and Follow-Your-Emoji (FYE) (Ma et al., 2024) are trained on much larger-scale datasets and are not suitable for a direct comparison.
Method Training Dataset FID $\downarrow$ CSIM $\uparrow$ ARD $\downarrow$
AniPortrait (Wei et al., 2024) VFHQ (Xie et al., 2022), CelebV-HQ (Zhu et al., 2022) 66.61 0.7226 2.9146
FYE (Ma et al., 2024) HDTF (Zhang et al., 2021), VFHQ (Xie et al., 2022), their collected dataset (Ma et al., 2024) 60.05 0.7558 3.0822
FOMM (Siarohin et al., 2019b) VoxCeleb1 (Nagrani et al., 2017) 80.00 0.6010 1.8331
DaGAN (Hong et al., 2022) VoxCeleb1 (Nagrani et al., 2017) 85.32 0.5743 2.0604
MCNet (Hong & Xu, 2023) VoxCeleb1 (Nagrani et al., 2017) 82.72 0.5618 1.6970
MRFA (Tao et al., 2024) VoxCeleb1 (Nagrani et al., 2017) 77.63 0.5962 1.5903
Ours VoxCeleb1 (Nagrani et al., 2017) 76.47 0.6142 1.6234
We also present a quantitative comparison with state-of-the-art methods for cross-identity reenactment in Tab. 3. We use FID (Heusel et al., 2017) for image quality evaluation, Average Rotation Distance (ARD) for motion transfer evaluation following Tao et al. (2024), and cosine similarity (CSIM) for identity preservation following Ha et al. (2020). Diffusion-based AniPortrait (Wei et al., 2024) and Follow-Your-Emoji (FYE) (Ma et al., 2024) are trained on much larger-scale datasets and are excluded from the comparison. Our method performs the best in terms of image quality and identity preservation, and is comparable in terms of motion transfer, confirming the effectiveness of our approach. Although MRFA (Tao et al., 2024) can transfer motion well, its output frame quality may be low and it may not preserve source identity effectively. Although diffusion-based methods can achieve even better image quality and identity preservation performance, they struggle to transfer motion faithfully, which leads to undesirable visual quality.
Table 4: Ablation study on the code allocation scheme. We present the results of different codebook splitting schemes.
Method FID $\downarrow$ PSNR $\uparrow$ $\mathcal{L}_{1}$downarrow LPIPS $\downarrow$ AKD $\downarrow$ AED $\downarrow$
Sharing all codes 43.23 25.12 0.0359 0.1860 1.2124 0.1065
Splitting the codes equally 42.52 25.20 0.0358 0.1857 1.1893 0.1075
Code Allocation (Ours) 43.15 25.30 0.0355 0.1846 1.2039 0.1071
Effect of the code allocation scheme. We propose a novel code allocation scheme for motion and appearance codebooks that assigns different codes to corresponding scales. This allows certain codes to be shared across multiple scales, facilitating the transfer of information between them. To assess the effectiveness of our code allocation scheme, we conduct an ablation study, as shown in Tab. 4. We compare two alternative codebook splitting schemes: sharing all codes across all scales and splitting the codes equally among the scales. As demonstrated in Tab. 4, our code allocation scheme generally achieves the best overall performance. These results highlight the effectiveness and superiority of our proposed code allocation scheme.
A.3 Limitations
A limitation of our method is the identity leakage problem in cross-identity reenactment, where the face in the generated video tends to have a shape similar to that of the driving face rather than the source face. This issue arises from the keypoint-based motion flow estimator that we adopt to produce the initial coarse motion flow and the driving keypoints for multi-scale motion codebook compensation. Although this motion flow estimator is robust to non-facial motion, such as hair and neck movement, by learning unsupervised keypoints on talking heads, the keypoints also inherently model facial shapes, which leads to the entanglement of motion and shape. Relative motion transfer (Siarohin et al., 2019b) can alleviate this problem and is widely adopted by previous methods (Hong et al., 2022; Hong & Xu, 2023; Tao et al., 2024).